OpenStudy (anonymous):

what is the derivative of 2x^2(1-x^2)?

OpenStudy (diyadiya):

\[2x^2(1-x^2)=2x^2-2x^4\]

OpenStudy (anonymous):

You can use the product rule here, but I think it's easier to distribute, and use the rule for polynomial terms .

OpenStudy (anonymous):

As Diyadiya did above....

OpenStudy (diyadiya):

\[\frac{d}{dx}2x^2-\frac{d}{dx}2x^4\]

OpenStudy (anonymous):

gotcha! than you!

OpenStudy (anonymous):

on the same problem. how do i find relative extrema using 2nd derivative?

OpenStudy (anonymous):

Extrema can be identified by the first derivative equal to zero. If the second derivative is positive at this value of x, the extrema is a minimum; if negative, a maximum.